# Classification of locality preserving symmetries on spin chains

### v1

@@ Appendix A Introduction
Non-trivial locality preserving automorphisms appear in various guises in the study of topological phases of strongly interacting quantum matter, and often serve to characterize and even classify the phases under investigation [3]. Examples include the appearance at stroboscopic times of the shift on the boundary of many-body localized Floquet insulators [4, 5, 6], the exotic symmetries appearing on the boundaries of topological matter [7, 8], and the equivariant automorphisms that can entangle symmetry protected trivial (SPT) phases [9, 6]. These connections have motivated a growing body of work that aims to understand the topological phases of locality preserving automorphisms, possibly in the presence of symmetry [2, 10, 11, 12, 13, 14, 15, 16].
In this paper we study representations of finite groups $G$ by locality preserving automorphisms on spin chains. Such locality preserving symmetries arise at the boundaries of two-dimensional SPTs whose bulk invariant manifests itself as an $H^{3}(G,U(1))$-valued anomaly of the boundary symmetry [17]. We prove the folk knowledge that this anomaly classifies locality preserving symmetries on spin chains up to stable equivalence. That is, up to conjugation by finite depth quantum circuits and stacking with factorized group actions.
@@
As a corollary we show that the obstruction to covariant right restrictions is given by the inverse of the slant product of the anomaly. Having a good handle on this quantity is significant because it plays an important role in characterizing the anyon content of the gauged bulk SPT corresponding to the boundary symmetry under consideration [18, 19]. Note in particular that any symmetry which admits covariant right restrictions can be presented by finite depth quantum circuits with covariant gates, a highly non-trivial property.
The paper is structured as follows. In Section B we introduce locality preserving symmetries on spin chains and stable equivalence between them, and state our main Theorem. We define the anomaly in Section C and state its basic properties. In Section D we construct for each element of $H^{3}(G,U(1))$ an explicit symmetry with that element as its anomaly. In Section E we define the obstruction to covariant right restrictions and show that it vanishes for symmetries with trivial anomaly. This fact is then used in Section F to prove the main Theorem. Appendix G collects basic definitions of group cohomology. Basic properties of the anomaly and of the obstruction to covariant right restrictions are proved in Appendices H and I respectively. In Appendix J we show that the obstruction to covariant right restrictions is given by the slant product of the anomaly. Finally, Appendix K presents an example which shows that stable equivalence is needed in order for the classification by the anomaly to hold.
@@ B.1 Spin chains, quantum cellular automata, and finite depth quantum circuits
A spin chain is an assignment of a full matrix algebra ${\mathcal{A}}_{j}\simeq\operatorname{\mathrm{End}}({\mathbb{C}}^{d_{j}})$ to each site $j\in{\mathbb{Z}}$ of the chain with uniformly bounded dimensions $d_{j}$. We write ${\mathcal{A}}$ for the corresponding algebra of quasi-local observables, and for any $\Gamma\subset{\mathbb{Z}}$ we write ${\mathcal{A}}_{\Gamma}$ for the $C^{*}$-algebra of quasi-local observables supported in $\Gamma$. We will write ${\mathcal{A}}_{\geq j}$ for ${\mathcal{A}}_{[j,\infty)}$ and ${\mathcal{A}}_{<j}$ for ${\mathcal{A}}_{(-\infty,j-1]}$. We will usually refer to the quasi-local algebra ${\mathcal{A}}$ itself as the ‘spin chain’, it being understood that there is a fixed preferred assignment of on-site algebras $j\mapsto{\mathcal{A}}_{j}\subset{\mathcal{A}}$ to sites of ${\mathbb{Z}}$.
For any $\Gamma\subset{\mathbb{Z}}$ we write $\Gamma^{(r)}:=\{j\in{\mathbb{Z}}\,:\,\operatorname{\mathrm{dist}}(j,\Gamma)\leq r\}$ for the ‘$r$-fattening’ of $\Gamma$.
A quantum cellular automaton (QCA) on a spin chain ${\mathcal{A}}$ is a *-automorphism $\alpha:{\mathcal{A}}\rightarrow{\mathcal{A}}$ for which there exists $r\geq 0$ such that $\alpha(x)\in{\mathcal{A}}_{\mathrm{supp}\,(x)^{(r)}}$ for all $x\in{\mathcal{A}}$.
The range of a QCA is the smallest $r$ for which this holds. The inverse of a QCA of range $r$ is also a QCA of range $r$ ([10, Lemma 3.1]), a fact which we will use without comment throughout the paper. The quantum cellular automata on ${\mathcal{A}}$ form a subgroup of $\operatorname{\mathsf{Aut}}({\mathcal{A}})$ which we denote by $\operatorname{\mathsf{QCA}}({\mathcal{A}})$.
Let $\{I_{a}\}_{a\in{\mathbb{Z}}}$ be a partition of ${\mathbb{Z}}$ into intervals $I_{a}\subset{\mathbb{Z}}$ of bounded size. Suppose we have for each $a\in{\mathbb{Z}}$ a unitary $U_{a}\in{\mathcal{A}}_{I_{a}}$, then we can define a QCA $\beta$ which is given by conjugation with the (formal) unitary $\bigotimes_{a\in{\mathbb{Z}}}U_{a}$. Any QCA of this form is called a block partitioned QCA. The intervals $I_{a}$ are called the blocks of the block partitioned QCA, and $\left|I_{a}\right|$ is the size of block $I_{a}$. The unitaries $U_{a}$ are called gates. The composition of $n$ block partitioned QCAs is called a depth $n$ quantum circuit, or simply a finite depth quantum circuit (FDQC).
@@ B.2 Locality preserving symmetries
Let $G$ be a finite group which will be fixed throughout the paper. We write $\bar{g}=g^{-1}$ for the inverse of any group element $g\in G$. A locality preserving symmetry on ${\mathcal{A}}$ is a group homomorphism $\alpha:G\rightarrow\operatorname{\mathsf{QCA}}({\mathcal{A}})$. That is, for each $g\in G$ we have a quantum cellular automaton $\alpha^{(g)}$ such that $\alpha^{(1)}=\mathrm{id}$ and $\alpha^{(g)}\circ\alpha^{(h)}=\alpha^{(gh)}$ for all $g,h\in G$. The range of a locality preserving symmetry is the largest range of its component QCAs. We say a locality preserving symmetry $\alpha$ is decoupled if there is a partition $\{I_{a}\}_{a\in{\mathbb{Z}}}$ of ${\mathbb{Z}}$ into intervals $I_{a}$ of bounded size such that $\alpha=\bigotimes_{a\in{\mathbb{Z}}}\,\alpha_{a}$ for symmetries $\alpha_{a}$ supported on $I_{a}$. In particular, any symmetry of range 0 is decoupled.
In the rest of this work we will refer to locality preserving symmetries simply as symmetries. If we want to specify the group $G$ then we speak of $G$-symmetries. We denote the set of all $G$-symmetries by $\operatorname{\mathsf{Sym}}_{G}$.
@@ B.3 Equivalence and stable equivalence
Two symmetries $\alpha$ and $\alpha^{\prime}$ are stably equivalent, denoted by $\alpha\sim\alpha^{\prime}$, if there exists decoupled symmetries $\beta$ and $\beta^{\prime}$ such that $\alpha\otimes\beta\sim_{0}\alpha^{\prime}\otimes\beta^{\prime}$. Stable equivalence is an equivalence relation on $\operatorname{\mathsf{Sym}}_{G}$, and $(\operatorname{\mathsf{Sym}}_{G}/\sim)$ is an abelian monoid with multiplication induced by stacking. (We will show later that it is in fact a group, i.e. there are inverses.) Note that any decoupled symmetry is stably equivalent to a symmetry of range zero.
@@ B.4 Main result
The monoid $(\operatorname{\mathsf{Sym}}_{G}/\sim)$ is in fact a group. There is a map $\Omega:\operatorname{\mathsf{Sym}}_{G}\rightarrow H^{3}(G,U(1))$ which assigns to each symmetry $\alpha$ its anomaly and which lifts to an isomorphism of groups $(\operatorname{\mathsf{Sym}}_{G}/\sim)\cong H^{3}(G,U(1))$.
In particular, two $G$-symmetries $\alpha$ and $\beta$ are stably equivalent if and only if their anomalies are equal:
@@
In Appendix K we describe a symmetry $\alpha$ with trivial anomaly $\Omega(\alpha)=[1]$ which is nevertheless not equivalent to a decoupled symmetry. This shows that the notion of stable equivalence is indeed necessary for the classification by the anomaly to hold.
@@ Appendix C The anomaly of a locality preserving symmetry
Given a right restriction $\alpha_{\geq j}$ of defect size $L$ there are local unitaries $\operatorname{\Phi}_{j}(g,h)\in{\mathcal{A}}_{[j-L,j+L+R]}$, called fusion operators associated to $\alpha_{\geq j}$, such that
@@
$1=\frac{\omega_{j}(g,h,k)\omega_{j}(f,gh,k)\omega_{j}(f,g,h)}{\omega_{j}(fg,h,k)\omega_{j}(f,g,hk)}$
and the corresponding group cohomology class $[\omega_{j}]\in H^{3}(G,U(1))$ depends only on the symmetry $\alpha$. We thus obtain a well defined map
@@
which we call the anomaly. If $\operatorname{\Omega}(\alpha)=[1]$ is the identity element of $H^{3}(G,U(1))$ then we say that $\alpha$ has trivial anomaly.
@@
Proof.
Can be found in Appendix H. ∎
@@ Appendix D Examples
Consider the spin chain with local degrees of freedom ${\mathcal{A}}_{x}\simeq\operatorname{\mathrm{End}}\big{(}{\mathbb{C}}^{\left|G\right|}\big{)}$. Define unitaries $U_{j,j+1}^{(g)}\in{\mathcal{A}}_{\{j,j+1\}}$ by
$U_{j,j+1}^{(g)}|g_{j},g_{j+1}\rangle=\omega(g,g_{j+1},\bar{g}_{j+1}g_{j})|g_{j},g_{j+1}\rangle.$
Note that the $U_{j,j+1}^{(g)}$ commute with each other for all $j\in{\mathbb{Z}}$ and for all $g\in G$.
Define $\alpha^{(g)}$ as the composition $\alpha_{3}^{(g)}\circ\alpha_{2}^{(g)}\circ\alpha_{1}^{(g)}$ of three block partitioned QCAs. The blocks of $\alpha_{1}^{(g)}$ are neighbouring pairs of sites $\{2a,2a+1\}$ and the corresponding gates are $U^{(g)}_{2a,2a+1}$. Similarly, $\alpha_{2}^{(g)}$ has blocks $\{2a-1,2a\}$ and corresponding gates $U_{2a-1,2a}^{(g)}$. Finally, the block partitioned QCA $\alpha_{3}^{(g)}$ has the singletons $\{a\}$ as blocks with the left action $L^{(g)}|h\rangle=|gh\rangle$ as gates. See Figure 1.
@@
Let $I=[a,b]\subset{\mathbb{Z}}$ be a finite interval and let $\alpha_{I}^{(g)}$ be the FDQC obtained from $\alpha^{(g)}$ by only retaining the gates that are supported on $I$, see Figure 2 The product of the finite number of gates of $\alpha_{I}^{(g)}$ then defines a unitary $U_{I}^{(g)}$ so that $\alpha_{I}^{(g)}=\mathrm{Ad}[U_{I}^{(g)}]$. Note that $\alpha^{(g)}=\lim_{a\uparrow\infty}\alpha_{[-a,a]}^{(g)}$ in the strong topology.
@@
It follows immediately from this lemma that $\alpha^{(g)}\alpha^{(h)}=\alpha^{(gh)}$ for all $g,h\in G$, so that $g\mapsto\alpha^{(g)}$ is indeed a group homomorphism.
Let $\alpha_{\geq}^{(g)}:=\lim_{b\uparrow\infty}\alpha_{[0,b]}^{(g)}$. Then $\alpha_{\geq}$ is a right restriction of $\alpha$ and
@@ E.1 Right restrictions that are group homomorphisms
The following Lemma says that any symmetry with trivial anomaly admits right restrictions that are group homomorphisms, at least after stacking with an auxiliary gauge degree of freedom. The proof is closely analogous to [22, Theorem 3.1.6].
@@
Let $\alpha$ be a symmetry of range $R$ with trivial anomaly. For any site $j\in{\mathbb{Z}}$ we can enlarge the on-site algebra ${\mathcal{A}}_{j}$ by tensoring with a gauge degree of freedom $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$, obtaining an enlarged spin chain $\widetilde{\mathcal{A}}$. Then the symmetry $\tilde{\alpha}=\alpha\otimes\mathrm{id}$ on $\widetilde{\mathcal{A}}$ admits a right restriction $\tilde{\beta}_{\geq j}$ at site $j$ of defect size $5R$ such that $g\mapsto\tilde{\beta}_{\geq j}^{(g)}$ is a group homomorphism.
@@
Let $\alpha_{\geq j}$ be a right restriction of $\alpha$ at $j\in{\mathbb{Z}}$ with defect size $2R$. We drop $j$ from the notation in the remainder of this proof. Let $\operatorname{\Phi}(g,h)\in{\mathcal{A}}_{[j-2R,j+3R]}$ be the fusion operators associated to this right restriction. Since $\alpha$ has trivial anomaly, by Eq. C.1 we can choose the phases of the fusion operators such that
@@
for all $f,g,h\in G$. Let $\tilde{\alpha}_{\geq}=\alpha_{\geq}\otimes\mathrm{id}$, which is a left restriction for $\tilde{\alpha}=\alpha\otimes\mathrm{id}$ of defect size $2R$. Define unitaries

### v2

@@ Appendix A Introduction
Non-trivial locality preserving automorphisms appear in various guises in the study of topological phases of strongly interacting quantum matter, and often serve to characterize and even classify the phases under investigation [3]. Examples include the appearance at stroboscopic times of the shift on the boundary of many-body localized Floquet insulators [4, 5, 6], the exotic symmetries appearing on the boundaries of topological matter [7, 8, 9], and the equivariant automorphisms that can entangle symmetry protected trivial (SPT) phases [10, 6]. These connections have motivated a growing body of work that aims to understand the topological phases of locality preserving automorphisms, possibly in the presence of symmetry [2, 11, 12, 13, 14, 15, 16, 17].
In this paper we study representations of finite groups $G$ by locality preserving automorphisms on spin chains [18, 19]. Such representations can be regarded as the group case of categorical symmetries on spin chains [20, 21, 22]. They also arise at the boundaries of two-dimensional SPTs, whose bulk invariant manifests itself as an $H^{3}(G,U(1))$-valued anomaly of the boundary symmetry [23]. We prove the folk knowledge that this anomaly classifies locality preserving symmetries on spin chains up to stable equivalence. That is, up to conjugation by finite depth quantum circuits and stacking with factorized group actions.
@@
As a corollary, we show that the obstruction to covariant right restrictions is given by the inverse of the slant product of the anomaly. Having a good handle on this quantity is significant because it plays an important role in characterizing the anyon content of the gauged bulk SPT corresponding to the boundary symmetry under consideration [24, 25]. Note in particular that any symmetry which admits covariant right restrictions can be presented by finite depth quantum circuits with covariant gates, a highly non-trivial property.
The paper is structured as follows. In Section B, we introduce locality preserving symmetries on spin chains and stable equivalence between them, and state our main Theorem. We define the anomaly in Section C, and state its basic properties. In Section D, we construct for each element of $H^{3}(G,U(1))$ an explicit symmetry with that element as its anomaly. In Section E, we define the obstruction to covariant right restrictions and show that it vanishes for symmetries with trivial anomaly. This fact is then used in Section F to prove the main Theorem. Appendix G collects basic definitions of group cohomology. Basic properties of the anomaly and of the obstruction to covariant right restrictions are proved in Appendices H and I respectively. In Appendix K we show that the obstruction to covariant right restrictions is given by the slant product of the anomaly. Finally, Appendix L presents an example which shows that stable equivalence is needed in order for the classification by the anomaly to hold.
@@ B.1 Spin chains, quantum cellular automata, and finite depth quantum circuits
A spin chain $C^{*}$-algebra ${\mathcal{A}}$ is defined in the standard way, that we recall now. To any site $j\in{\mathbb{Z}}$, we associate an $d_{j}$-dimensional on-site Hilbert space ${\mathbb{C}}^{d_{j}}$, with associated matrix algebra ${\mathcal{A}}_{j}\simeq\operatorname{\mathrm{End}}({\mathbb{C}}^{d_{j}})$. We assume that there is a $d_{\max}$ such that $d_{j}\leq d_{\max}$. The algebra ${\mathcal{A}}_{j}\simeq\operatorname{\mathrm{End}}({\mathbb{C}}^{d_{j}})$ is equipped with its natural operator norm and *-operation (Hermitian adjoint of a matrix) making it into a $C^{*}$-algebra. The spin chain algebra ${\mathcal{A}}$ is the inductive limit of algebras ${\mathcal{A}}_{S}=\otimes_{j\in S}{\mathcal{A}}_{j}$, with $S$ a finite subset of ${\mathbb{Z}}$. It comes naturally equipped with local subalgebras ${\mathcal{A}}_{X},X\subset{\mathbb{Z}}$. We refer to standard references [27, 28, 29, 30, 31] for more background and details. We will write ${\mathcal{A}}_{\geq j}$ for ${\mathcal{A}}_{[j,\infty)}$ and ${\mathcal{A}}_{<j}$ for ${\mathcal{A}}_{(-\infty,j-1]}$. We will usually refer to the quasi-local algebra ${\mathcal{A}}$ itself as the spin chain, it being understood that there is a fixed preferred assignment of on-site algebras $j\mapsto{\mathcal{A}}_{j}\subset{\mathcal{A}}$ to sites of ${\mathbb{Z}}$.
For any $\Gamma\subset{\mathbb{Z}}$ we write $\Gamma^{(r)}:=\{j\in{\mathbb{Z}}\,:\,\operatorname{\mathrm{dist}}(j,\Gamma)\leq r\}$ for the $r$-fattening of $\Gamma$. A quantum cellular automaton (QCA) on a spin chain ${\mathcal{A}}$ is a *-automorphism $\alpha:{\mathcal{A}}\rightarrow{\mathcal{A}}$ for which there exists $r\geq 0$ such that $\alpha({\mathcal{A}}_{X})\subset{\mathcal{A}}_{X^{(r)}}$ for any $X\subset{\mathbb{Z}}$.
The range of a QCA is the smallest $r$ for which this holds. The inverse of a QCA of range $r$ is also a QCA of range $r$ ([11, Lemma 3.1]), a fact which we will use without comment throughout the paper. The quantum cellular automata on ${\mathcal{A}}$ form a subgroup of $\operatorname{\mathsf{Aut}}({\mathcal{A}})$ which we denote by $\operatorname{\mathsf{QCA}}({\mathcal{A}})$.
Let $\{I_{a}\}_{a\in{\mathbb{Z}}}$ be a partition of ${\mathbb{Z}}$ into intervals $I_{a}\subset{\mathbb{Z}}$ of bounded size. Suppose we have for each $a\in{\mathbb{Z}}$ a unitary $U_{a}\in{\mathcal{A}}_{I_{a}}$, then we can define a QCA $\beta$ by the formal infinite product
$\beta=\bigotimes_{a\in{\mathbb{Z}}}\mathrm{Ad}(U_{a}).$
This yields a well-defined automorphism, as one can first define its action on ${\mathcal{A}}_{X}$ with finite $X$ and then extend by density. Any QCA of this form is called a block partitioned QCA. The intervals $I_{a}$ are called the blocks of the block partitioned QCA, and $\left|I_{a}\right|$ is the size of block $I_{a}$. The unitaries $U_{a}$ are called gates. The composition of $n$ block partitioned QCAs is called a depth $n$ quantum circuit, or simply a finite depth quantum circuit (FDQC).
@@ B.2 Locality preserving symmetries
Let $G$ be a finite group which will be fixed throughout the paper. We write $\bar{g}=g^{-1}$ for the inverse of any group element $g\in G$. A locality preserving symmetry on ${\mathcal{A}}$ is a group homomorphism $\alpha:G\rightarrow\operatorname{\mathsf{QCA}}({\mathcal{A}})$. That is, for each $g\in G$ we have a quantum cellular automaton $\alpha^{(g)}$ such that $\alpha^{(1)}=\mathrm{id}$ and $\alpha^{(g)}\circ\alpha^{(h)}=\alpha^{(gh)}$ for all $g,h\in G$. The range of a locality preserving symmetry is the largest range of its component QCAs. We say a locality preserving symmetry $\alpha$ is decoupled iff. every $\alpha^{(g)},g\in G$ is a block-partitioned QCA (as defined above) where the blocks can be chosen to be $g$-independent. Alternatively, this means that we can write formally $\alpha=\bigotimes_{a\in{\mathbb{Z}}}\,\alpha_{a}$ for symmetries $\alpha_{a}$ supported on the blocks $I_{a}$. In particular, any symmetry of range 0 is decoupled.
In the rest of this work we will refer to locality preserving symmetries simply as symmetries. If we want to specify the group $G$ then we speak of $G$-symmetries. We denote the set of all $G$-symmetries on arbitrary spin chains by $\operatorname{\mathsf{Sym}}_{G}$.
@@ B.3 Equivalence and stable equivalence
Two symmetries $\alpha$ and $\alpha^{\prime}$ are stably equivalent, denoted by $\alpha\sim\alpha^{\prime}$, if there exists symmetries $\beta$ and $\beta^{\prime}$ of range 00 such that $\alpha\otimes\beta\sim_{0}\alpha^{\prime}\otimes\beta^{\prime}$. Stable equivalence is an equivalence relation on $\operatorname{\mathsf{Sym}}_{G}$, and $(\operatorname{\mathsf{Sym}}_{G}/\sim)$ is an abelian monoid with multiplication induced by stacking. (We will show later that it is in fact a group, i.e. there are inverses.) It is easy to check that any decoupled symmetry is stably equivalent to a symmetry of range zero. This implies also that $\alpha,\alpha^{\prime}$ are stably equivalent whenever there exist decoupled symmetries $\beta,\beta^{\prime}$ such that $\alpha\otimes\beta\sim_{0}\alpha^{\prime}\otimes\beta^{\prime}$. This fact will be used throughout the paper without further mention.
@@ B.4 Main result
The monoid $(\operatorname{\mathsf{Sym}}_{G}/\sim)$ is in fact a group. There is a map $\Omega:\operatorname{\mathsf{Sym}}_{G}\rightarrow H^{3}(G,U(1))$ which assigns to each symmetry $\alpha$ a 3-cohomology class, which we will call its anomaly, and which lifts to an isomorphism of groups $(\operatorname{\mathsf{Sym}}_{G}/\sim)\cong H^{3}(G,U(1))$.
In particular, two $G$-symmetries $\alpha$ and $\beta$ are stably equivalent if, and only if, their anomalies are equal:
@@
In Appendix L we describe a symmetry $\alpha$ with trivial anomaly $\Omega(\alpha)=[1]$ which is nevertheless not equivalent to a decoupled symmetry. This shows that the notion of stable equivalence is indeed necessary for the classification by the anomaly to hold.
@@ Appendix C The anomaly of a locality preserving symmetry
Given a right restriction $\alpha_{\geq j}$ of defect size $L$, there are local unitaries $\operatorname{\Phi}_{j}(g,h)\in{\mathcal{A}}_{[j-L,j+L+R]}$, called fusion operators associated to $\alpha_{\geq j}$, such that
@@
$1=\frac{\omega_{j}(g,h,k)\omega_{j}(f,gh,k)\omega_{j}(f,g,h)}{\omega_{j}(fg,h,k)\omega_{j}(f,g,hk)},$
and the corresponding group cohomology class $[\omega_{j}]\in H^{3}(G,U(1))$ depends only on the symmetry $\alpha$, i.e. the cohomology class is independent of the site $j$ and the choice of right restriction $\alpha_{\geq j}$. We thus obtain a well defined map
@@
which we call the anomaly. If $\operatorname{\Omega}(\alpha)=[1]$ is the identity element of $H^{3}(G,U(1))$, then we say that $\alpha$ has trivial anomaly.
@@
The proof can be found in Appendix H.
@@ Appendix D Examples
Consider the spin chain with on-site algebras ${\mathcal{A}}_{x}\simeq\operatorname{\mathrm{End}}\big{(}{\mathbb{C}}^{\left|G\right|}\big{)}$. Define unitaries $V_{j,j+1}^{(g)}\in{\mathcal{A}}_{\{j,j+1\}}$ by
$V_{j,j+1}^{(g)}|g_{j},g_{j+1}\rangle=\omega(g,g_{j+1},\bar{g}_{j+1}g_{j})|g_{j},g_{j+1}\rangle.$
Note that the $V_{j,j+1}^{(g)}$ commute with each other for all $j\in{\mathbb{Z}}$ and for all $g\in G$.
Define $\alpha^{(g)}$ as the composition $\alpha_{3}^{(g)}\circ\alpha_{2}^{(g)}\circ\alpha_{1}^{(g)}$ of three block partitioned QCAs. The blocks of $\alpha_{1}^{(g)}$ are neighbouring pairs of sites $\{2a,2a+1\}$ and the corresponding gates are $V^{(g)}_{2a,2a+1}$. Similarly, $\alpha_{2}^{(g)}$ has blocks $\{2a-1,2a\}$ and corresponding gates $V_{2a-1,2a}^{(g)}$. Finally, the block partitioned QCA $\alpha_{3}^{(g)}$ has the singletons $\{a\}$ as blocks with the left action $L^{(g)}|h\rangle=|gh\rangle$ as gates. See Figure 1.
@@
Let $I=[a,b]\subset{\mathbb{Z}}$ be a finite interval and let $\alpha_{I}^{(g)}$ be the FDQC obtained from $\alpha^{(g)}$ by only retaining the gates that are supported on $I$, see Figure 2. The product of the finite number of gates of $\alpha_{I}^{(g)}$ then defines a unitary $U_{I}^{(g)}$ so that $\alpha_{I}^{(g)}=\mathrm{Ad}[U_{I}^{(g)}]$. Note that $\alpha^{(g)}=\lim_{a\uparrow\infty}\alpha_{[-a,a]}^{(g)}$ in the strong topology.
@@
It follows immediately from this lemma that $\alpha^{(g)}\alpha^{(h)}=\alpha^{(gh)}$ for all $g,h\in G$, so that $g\mapsto\alpha^{(g)}$ is indeed a group homomorphism. Let $\alpha_{\geq}^{(g)}:=\lim_{b\uparrow\infty}\alpha_{[0,b]}^{(g)}$. Then $\alpha_{\geq}$ is a right restriction of $\alpha$ and
@@ E.1 Right restrictions that are group homomorphisms
The following Lemma says that any symmetry with trivial anomaly admits right restrictions that are group homomorphisms, at least after stacking with a local degree of freedom. The proof is closely analogous to [33, Theorem 3.1.6].
A local degree of freedom at site $j$ is a spin chain where the on-site algebras ${\mathcal{A}}_{i}\simeq{\mathbb{C}}$ are trivial for $i\neq j$. For convenience, also the only non-trivial on-site algebra ${\mathcal{A}}_{j}$ is sometimes also called the local degree of freedom. The procedure of stacking a local degree of freedom ${\mathcal{A}}_{j}^{\prime}$ with a spin chain ${\mathcal{A}}$ results in a spin chain that is isomorphic to ${\mathcal{A}}\otimes{\mathcal{A}}^{\prime}_{j}$. A simultaneous stacking of a countable number of (uniformly upper bounded) local degrees of freedom at different sites with a spin chain ${\mathcal{A}}$ still produces a spin chain with uniformly upper bounded on-site dimensions.
@@
Let $\alpha$ be a symmetry of range $R$ with trivial anomaly. For any site $j\in{\mathbb{Z}}$ we can stack the on-site algebra ${\mathcal{A}}_{j}$ by a local degree of freedom $\operatorname{\mathrm{End}}({\mathbb{C}}^{\left|G\right|})$, obtaining an enlarged spin chain $\widetilde{\mathcal{A}}$. Then the composed symmetry $\tilde{\alpha}=\alpha\otimes\mathrm{id}$ on $\widetilde{\mathcal{A}}$ admits a right restriction $\tilde{\beta}_{\geq j}$ at site $j$ of defect size $5R$ such that $g\mapsto\tilde{\beta}_{\geq j}^{(g)}$ is a group homomorphism.
@@
Let $\alpha_{\geq j}$ be a right restriction of $\alpha$ at $j\in{\mathbb{Z}}$ with defect size $2R$. We drop $j$ from the notation in the remainder of this proof. Let $\operatorname{\Phi}(g,h)\in{\mathcal{A}}_{[j-2R,j+3R]}$ be the fusion operators associated to this right restriction. Since $\alpha$ has trivial anomaly, by Eq. (C.1) we can choose the phases of the fusion operators such that
@@
for all $f,g,h\in G$. Let $\tilde{\alpha}_{\geq}=\alpha_{\geq}\otimes\mathrm{id}$, which is a right restriction for $\tilde{\alpha}=\alpha\otimes\mathrm{id}$ of defect size $2R$. Define unitaries